# On Selection of Cross-Section Averages in Non-stationary Environments

### v3

@@ 4 Conclusion
This study is the first to examine selection of an optimal set of CAs in CCE and related settings by ICs inspired by Bai and Ng, (2002) when latent factors are non-stationary. In particular, we use mild integration in order to explore varying degrees of factor persistence, and demonstrate that ICs remain consistent without any modifications. However, the more persistent common factors are, the worse small sample performance becomes, and ICs regain selection consistency only in very large samples (i.e. $(N,T)>300$, according to our experiments). Importantly, a diverging penalty suggested by Bai, (2004) under non-stationary factors for PC renders ICs inconsistent in the case of CAs. Therefore, our recommendation for CCE practitioners is not to take prescriptions from the PC literature automatically, and interpret IC results with caution in presence of highly persistent data, unless $N,T$ are substantial.
@@ 5.1 Assumptions
Assumption 1.
  • (a)
    $\{\mathbf{f}_{t}\}$ is a mildly integrated process as defined in Magdalinos and Phillips, (2009). In particular,
    $\mathbf{f}_{t}=\mathbf{R}_{f,T}\mathbf{f}_{t-1}+\mathbf{u}_{f,t},\quad\mathbf{R}_{f,T}=\mathbf{I}_{m}-\frac{\mathbf{G}}{T^{\tau}},\quad\mathbf{G}>0\hskip 5.69054pt(diagonal),$
@@
  • (b)
    If $\tau=0$, then $\{\mathbf{f}_{t}\}$ is a stationary process with summable autocovariances.
@@ 5.2 Auxiliary Results
5.3 Proof of Proposition 1
In the proof we focus on the IC of De Vos and Stauskas, (2024) ($j=DVS$). Therefore, we will have $K=k$ throughout. The argument for the IC of Margaritella and Westerlund, (2023) ($j=MW$) is very similar, and we only need to take into account that we are working with the usual CCE setup, where $\widehat{\boldsymbol{\beta}}=\boldsymbol{\beta}+o_{p}(1)$ under the full set of $k+1$ CAs. We will comment on this at the end of the proof. For notational simplicity, we will use $\mathbf{q}_{M}\in\mathbb{R}^{k\times g}$ without the subscript $\mathbf{x}$. Also, because IC of DVS covers Examples 2 and 3 (see the main text), we denote factors as $\mathbf{F}_{\mathbf{x}}$.
@@ 5.3.1 Case of $M\subset M_{0}$
$\left\|\mathrm{tr}\left[\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}T^{\tau}\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right]\right\|$ $\leq\left\|\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}\right\|T^{\tau}\left\|\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right\|$
@@ 5.3.3 Proof Changes for IC of Margaritella and Westerlund, (2023)
5.3.3.1  Homogeneous $\boldsymbol{\beta}$
@@
$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})-\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}}-\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})$
@@
5.3.3.2  Heterogeneous $\boldsymbol{\beta}$
@@
$\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})\right\|=\begin{cases}O_{p}(T^{\tau}),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ O_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.98)
@@
with $d_{NT}\to_{p}d>0$ in analogy to the analysis in (5.3.3.1). Now, we need to assume $T^{\tau}N^{-1/2}\to 0$ in addition to $\tau\in(0,0.5)$ to isolate the $o_{p}(T^{-\tau})$ terms. Then in the under-specification case, we have
@@
as expected.
@@ 5.4 Additional Monte Carlo Results
5.4.1 Independent Idiosyncratics
@@ 5.4.3 Penalty by Bai, (2004)
5.4.4 Correlated Idiosyncratics

### v4

@@ 4 Conclusion
This study is the first to examine the selection of an optimal set of CAs in CCE and related settings by ICs inspired by Bai and Ng, (2002) when latent factors are non-stationary. In particular, we use mild integration to explore varying degrees of factor persistence and demonstrate that ICs remain consistent without any modifications. However, the more persistent common factors are, the worse their small sample performance becomes, and ICs regain selection consistency only in very large samples (i.e. $(N,T)>300$, according to our experiments). Importantly, a divergent penalty suggested by Bai, (2004) to address non-stationary factors in PC makes our ICs perform even worse in the case of CAs. Therefore, our recommendation for CCE/CAs practitioners is not to automatically take PC literature prescriptions and interpret IC results with caution in the presence of highly persistent data, unless $N,T$ are substantial.
@@ 5.1 Assumptions
Assumption 1. $\{\mathbf{f}_{t}\}$ is a mildly integrated process as defined in Magdalinos and Phillips, (2009), such that
$\mathbf{f}_{t}=\mathbf{R}_{f,T}\mathbf{f}_{t-1}+\mathbf{u}_{f,t},\quad\mathbf{R}_{f,T}=\mathbf{I}_{m}-\mathbf{G}T^{-\tau},\quad\mathbf{G}=\mathrm{diag}(g_{1},\ldots,g_{m}),\hskip 5.69054ptg_{j}\in(0,2),$
@@ 5.2 Auxiliary Results
5.3 Proposition 1 and the Proof
$\textit{(a)}\hskip 5.69054pt\textit{Under}\hskip 2.84526ptM_{0}\subset M,\hskip 2.84526pt\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=O_{p}(C_{N,T}^{-2});$ (A.8)
$\textit{(b)}\hskip 5.69054pt\textit{Under}\hskip 2.84526ptM\subset M_{0},\hskip 2.84526pt\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\tau b\ln(T)+\sum_{j:\lambda>0}\ln\left(\lambda_{j}(\boldsymbol{\Omega}_{M.M_{0}}^{0})\right)+R_{N,T},$ (A.9)
where $\boldsymbol{\Omega}_{M.M_{0}}^{0}=\operatorname*{plim}_{(N,T)\to\infty}\boldsymbol{\Omega}_{M.M_{0}}$ (positive semi-definite, non-zero), $|R_{N,T}|=o_{p}(1)$ is the remainder, $\lambda_{j}(\mathbf{A})$ is the $j$-th eigenvalue of $\mathbf{A}$, while $b$ is a number of strictly positive eigenvalues.
Proof. In the proof, we focus on the IC of De Vos and Stauskas, (2024) ($j=DVS$). Therefore, we will have $K=k$ throughout. The argument for the IC of Margaritella and Westerlund, (2023) ($j=MW$) is very similar, and we only need to take into account that we are working with the usual CCE setup, where $\widehat{\boldsymbol{\beta}}=\boldsymbol{\beta}+o_{p}(1)$ under the full set of $k+1$ CAs. We will comment on this at the end of the proof. For notational simplicity, we will use $\mathbf{q}_{M}\in\mathbb{R}^{k\times g}$ without the subscript $\mathbf{x}$. Also, because IC of DVS covers Examples 2 and 3 (see the main text), we denote factors as $\mathbf{F}_{\mathbf{x}}$.
@@ 5.3.1 Case of $M\subset M_{0}$
$\left|\mathrm{tr}\left[\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}T^{\tau}\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right]\right|$ $\leq\left\|\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}\right\|T^{\tau}\left\|\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right\|$
@@ 5.3.3 Proof Changes for IC of Margaritella and Westerlund, (2023)
Homogeneous $\boldsymbol{\beta}$
@@
$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})$
$-\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}}-\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})$
@@
Heterogeneous $\boldsymbol{\beta}$
@@
$\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})\right\|=\begin{cases}O_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ O_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.100)
@@
with $d_{NT}\to_{p}d>0$ in analogy to the analysis in (5.3.3). Now, we need to assume $T^{\tau}N^{-1/2}\to 0$ in addition to $\tau\in(0,0.5)$ to isolate the $o_{p}(T^{-\tau})$ terms. Then in the under-specification case, we have
@@
as expected. Note that the restrictions on $\tau$ discussed here are just sufficient conditions, since we do not see such need in the simulationd exercises.
@@ 5.4 Additional Monte Carlo Results
5.4.1 Idiosyncratics Independent Over Time
@@ 5.4.3 Penalty by Bai, (2004)
5.4.4 Idiosyncratics Correlated Over Time
@@ 5.4.5 Eigenvalue Ratio
5.4.6 Mis-selection with Varying $\tau$
Figure 1 presents the share of mis-selected number of cross-sectional averages using $IC^{MW}_{1}$ and $IC^{DVS}_{1}$ and $N=T=100$ (”small sample”) and $N=T=500$ (”large sample”), where we increase $\tau$ in $0.01$ steps from 0 to $0.95$. As in Table 1 of the manuscript the idiosyncratics $\mathbf{v}_{i,t}$ and $\varepsilon_{i,t}$ are uncorrelated over time, but weakly correlated across units. For the small sample, the share of mis-selected number of CAs remains relatively flat and below $20\%$ until $\tau$ reaches levels of around $0.5$, which is the midpoint. Then, it increases to almost $100\%$ with swings showcasing that the criteria become more unstable as $\tau$ increases. As expected, the degree of non-stationarity is much less of a problem for the large sample, and we observe misselection and instability of the criteria only for a high degree of non-stationarity. To shed more light on the importance of the sample size, we present in Figure 2 the share of misselected number of CAs with varying number of time periods from 25 to 500, $\tau=0.9$ and $N=[100,500]$. The share of misselection remains high for $N=100$ for both ICs, but declines for $N=500$. Again, this presents further evidence of the consistency of the ICs, but large samples are required to use them effectively.
(a) $IC^{DVS}_{1}$
(b) $IC^{MW}_{1}$
(a) $IC^{DVS}_{1}$
(b) $IC^{MW}_{1}$